L^{A}TEX 2ε font selection

\begin{matrix}c \\ ◯ \end{matrix} Copyright 1995 1996, L^{A}TEX3 Project Team. All rights reserved.

10 June 1996

Contents

1 Introduction 2 1.1 L^{A}TEX 2ε fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Further information . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Text fonts 4 2.1 Text font attributes . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Selection commands . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Parameters for author commands . . . . . . . . . . . . . . . . . . 8 2.5 Special font declaration commands . . . . . . . . . . . . . . . . . 9

3 Math fonts 10 3.1 Math font attributes . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 Selection commands . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.3 Declaring math versions . . . . . . . . . . . . . . . . . . . . . . . 13 3.4 Declaring math alphabets . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Declaring symbol fonts . . . . . . . . . . . . . . . . . . . . . . . . 14 3.6 Declaring math symbols . . . . . . . . . . . . . . . . . . . . . . . 15 3.7 Declaring math sizes . . . . . . . . . . . . . . . . . . . . . . . . . 16

4 Font installation 17 4.1 Font definition files . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.2 Font definition file commands . . . . . . . . . . . . . . . . . . . . 17 4.3 Font file loading information . . . . . . . . . . . . . . . . . . . . . 19 4.4 Size functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5 Encodings 21

5.1 The fontenc package . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.2 Encoding definition files . . . . . . . . . . . . . . . . . . . . . . . 22 5.3 Default definitions . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.4 Encoding defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.5 Case changing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

6 Miscellanea 26 6.1 Font substitution . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 6.2 Preloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.3 Accented characters . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.4 Naming conventions . . . . . . . . . . . . . . . . . . . . . . . . . 28

7 If you need to know more … 29

1 Introduction

This document describes the new font selection features of the L^{A}TEX Document Preparation System. It is intended for package writers who want to write fontloading packages similar to times or latexsym.

This document is only a brief introduction to the new facilities and is intended for package writers who are familiar with TEX fonts and L^{A}TEX packages. It is neither a user-guide nor a reference manual for fonts in L^{A}TEX 2ε.

1.1 L^{A}TEX 2ε fonts

The most important difference between L^{A}TEX 2.09 and L^{A}TEX 2ε is the way that fonts are selected. In L^{A}TEX 2.09, the Computer Modern fonts were built into the L^{A}TEX format, and so customizing L^{A}TEX to use other fonts was a major effort.

In L^{A}TEX 2ε, very few fonts are built into the format, and there are commands to load new text and math fonts. Packages such as times or latexsym allow authors to access these fonts. This document describes how to write similar font-loading packages.

The L^{A}TEX 2ε font selection system was first released as the ‘New Font Selection Scheme’ (NFSS) in 1989, and then in release 2 in 1993. L^{A}TEX 2ε includes NFSS release 2 as standard.

1.2 Overview

This document contains an overview of the new font commands of L^{A}TEX.

Section 2 describes the commands for selecting fonts in classes and packages. It lists the five L^{A}TEX font attributes, and lists the commands for selecting fonts. It also describes how to customize the author commands such as \textrm and \textit to suit your document design.

Section 3 explains the commands for controlling L^{A}TEX math fonts. It describes how to specify new math fonts and new math symbols.

Section 4 explains how to install new fonts into L^{A}TEX. It shows how L^{A}TEX font attributes are turned into TEX font names, and how to specify your own fonts using font definition files.

Section 5 discusses text font encodings. It describes how to declare a new encoding and how to define commands, such as \AE or \", which have different definitions in different encodings, depending on whether ligatures, etc. are available in the encoding.

Section 6 covers font miscellanea. It describes how L^{A}TEX performs font substitution, how to customize fonts that are preloaded in the L^{A}TEX format, and the naming conventions used in L^{A}TEX font selection.

1.3 Further information

For a general introduction to L^{A}TEX, including the new features of L^{A}TEX 2ε, you should read L^{A}TEX: A Document Preparation System, Leslie Lamport, Addison Wesley, 2nd ed, 1994.

A more detailed description of the L^{A}TEX font selection scheme is to be found in The L^{A}TEX Companion, Goossens, Mittelbach and Samarin, Addison Wesley, 1994.

The L^{A}TEX font selection scheme is based on TEX, which is described by its developer in The TEXbook, Donald E. Knuth, Addison Wesley, 1986, revised in 1991 to include the features of TEX 3.

Sebastian Rahtz’s psnfss software contains the software for using a large number of Type 1 fonts (including the Adobe Laser Writer 35 and the Monotype CD-ROM fonts) in L^{A}TEX. It should be available from the same source as your copy of L^{A}TEX.

The psnfss software uses fonts generated by Alan Jeffrey’s fontinst software. This can convert fonts from Adobe Font Metric format into a format readable by L^{A}TEX, including the generation of the .fd files described in Section 4. The fontinst software should be available from the same source as your copy of L^{A}TEX.

Whenever practical, L^{A}TEX uses the font naming scheme described in Filenames for fonts, Karl Berry, TUGboat 11(4), 1990.

The class-writer’s guide L^{A}TEX 2ε for Class and Package Writers describes the new L^{A}TEX features for writers of document classes and packages and is kept in clsguide.tex. Configuring L^{A}TEX is covered by the guide Configuration options for L^{A}TEX 2ε in cfgguide.tex whilst the philosophy behind our policy on modifying L^{A}TEX is described in Modifying L^{A}TEX in modguide.tex.

We are gradually turning the source code for L^{A}TEX into a L^{A}TEX document L^{A}TEX: the program. This document includes an index of L^{A}TEX commands and can be typeset from source2e.tex.

For more information about TEX and L^{A}TEX, please contact your local TEX Users Group, or the international TEX Users Group. Here are some useful addresses.

TEX Users Group, 1850 Union Street #1637, San Francisco, CA 94123, USA Fax: +1 415 982 8559 Email: tug@tug.org UK TUG, 1 Eymore Close, Selly Oak, Birmingham B29 4LB, UK

     Fax: +44 121 476 2159 Email: uktug-enquiries@tex.ac.uk

2 Text fonts

This section describes the commands available to class and package writers for specifying and selecting fonts.

2.1 Text font attributes

Every text font in L^{A}TEX has five attributes:

encoding This specifies the order that characters appear in the font. The two most common text encodings used in L^{A}TEX are Knuth’s ‘TEX text’ encoding, and the ‘TEX text extended’ encoding developed by the TEX Users Group members during a TEX Conference at Cork in 1990 (hence its informal name ‘Cork encoding’).

family The name for a collection of fonts, usually grouped under a common name by the font foundry. For example, ‘Adobe Times’, ‘ITC Garamond’, and Knuth’s ‘Computer Modern Roman’ are all font families.

series How heavy or expanded a font is. For example, ‘medium weight’, ‘narrow’ and ‘bold extended’ are all series.

shape The form of the letters within a font family. For example, ‘italic’, ‘oblique’ and ‘upright’ (sometimes called ‘roman’) are all font shapes.

size The design size of the font, for example ‘10pt’.

The possible values for these attributes are given short acronyms by L^{A}TEX. The most common values for the font encoding are:

OT1 TEX text
T1 TEX extended text
OML TEX math italic
OMS TEX math symbols
OMX TEX math large symbols
U Unknown
L⟨xx⟩ A local encoding

The ‘local’ encodings are intended for font encodings which are only locally available, for example a font containing an organisation’s logo in various sizes.

There are far too many font families to list them all, but some common ones are:

cmr Computer Modern Roman
cmss Computer Modern Sans
cmtt Computer Modern Typewriter
cmm Computer Modern Math Italic
cmsy Computer Modern Math Symbols
cmex Computer Modern Math Extensions
ptm Adobe Times
phv Adobe Helvetica
pcr Adobe Courier

The most common values for the font series are:

m Medium
b Bold
bx Bold extended
sb Semi-bold
c Condensed

The most common values for the font shape are:

n Normal (that is ‘upright’ or ‘roman’)
it Italic
sl Slanted (or ‘oblique’)
sc Caps and small caps

The font size is specified as a dimension, for example 10pt or 1.5in or 3mm. These five parameters specify every L^{A}TEX font, for example:

L^{A}TEX specification  Font  TEX font name
OT1 cmr m n 10pt  Computer Modern Roman 10pt  cmr10
OT1cmssmsl12pt Computer Modern Sans Oblique 12pt  cmssi12
OML cmm mit10pt Computer Modern Math Italic 10pt  cmmi10
T1 ptm bit18pt  Adobe Times Bold Italic 18pt  ptmbq at 18pt

These five parameters are displayed whenever L^{A}TEX gives an overfull box warning, for example:

   Overfull \hbox (3.80855pt too wide) in paragraph at lines 314--318
   []\OT1/cmr/m/n/10 Normally [] and [] will be iden-ti-cal,

The author commands for fonts set the five attributes:

Author command Attribute Value in article class

      \textrm{..} or \rmfamily family cmr
      \textsf{..} or \sffamily family cmss
      \texttt{..} or \ttfamily family cmtt
      \textmd{..} or \mdseries  series  m
      \textbf{..} or \bfseries  series  bx
       \textup{..} or \upshape shape n
       \textit{..} or \itshape shape it
       \textsl{..} or \slshape shape sl
       \textsc{..} or \scshape shape sc
                        \tiny  size  5pt
                  \scriptsize  size  7pt
                \footnotesize  size  8pt
                       \small  size  9pt
                  \normalsize  size  10pt
                       \large  size  12pt
                       \Large  size  14.4pt
                       \LARGE  size  17.28pt
                        \huge  size  20.74pt
                        \Huge  size  24.88pt

The values used by these commands are determined by the document class, using the parameters defined in Section 2.4.

Note that there are no author commands for selecting new encodings. These should be provided by packages, such as the fontenc package.

This section does not explain how L^{A}TEX font specifications are turned into TEX font names. This is described in Section 4.

2.2 Selection commands

The low-level commands used to select a text font are as follows.

\fontencoding {⟨encoding⟩}
\fontfamily {⟨family⟩}
\fontseries {⟨series⟩}
\fontshape {⟨shape⟩}

\fontsize {⟨size⟩} {⟨baselineskip⟩}

Each of these commands sets one of the font attributes; \fontsize also sets \baselineskip. The actual font in use is not altered by these commands, but the current attributes are used to determine which font to use after the next

\selectfont command.

\selectfont

Selects a text font, based on the current values of the font attributes.

Warning: there must be a \selectfont command immediately after any settings of the font parameters by (some of) the five \font⟨parameter⟩ commands, before any following text. For example, it is legal to say:

   \fontfamily{ptm}\fontseries{b}\selectfont Some text.

but it is not legal to say:

   \fontfamily{ptm} Some \fontseries{b}\selectfont text.

You may get unexpected results if you put text between a \font⟨parameter⟩ command and a \selectfont.

\usefont {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

A short hand for the equivalent \font… commands followed by a call to

\selectfont.

2.3 Internals

The current values of the font attributes are held in internal macros.

These hold the current values of the encoding, the family, the series, the shape, the size, the baseline skip, the main math size, the ‘script’ math size and the ‘scriptscript’ math size. The last three are accessible only within a formula; outside of math they may contain arbitrary values.

\f@encoding \f@family \f@series \f@shape \f@size \f@baselineskip \tf@size \sf@size \ssf@size

For example, to set the size to 12 without changing the baseline skip:

   \fontsize{12}{\f@baselineskip}

However, you should never alter the values of the internal commands directly; they must only be modified using the low-level commands like \fontfamily, \fontseries, etc. If you disobey this warning you might produce code that loops.

2.4 Parameters for author commands

The parameter values set by author commands such as \textrm and \rmfamily, etc. are not hard-wired into L^{A}TEX; instead these commands use the values of a number of parameters set by the document class and packages. For example, \rmdefault is the name of the default family selected by \textrm and \rmfamily. Thus to set a document in Adobe Times, Helvetica and Courier, the document designer specifies:

   \renewcommand{\rmdefault}{ptm}
   \renewcommand{\sfdefault}{phv}
   \renewcommand{\ttdefault}{pcr}
\encodingdefault
\familydefault
\seriesdefault
\shapedefault

The encoding, family, series and shape of the main body font. By default these are OT1, \rmdefault, m and n. Note that since the default family is \rmdefault, this means that changing \rmdefault will change the main body font of the document.

\rmdefault \sfdefault \ttdefault

The families selected by \textrm, \rmfamily, \textsf, \sffamily, \texttt

and \ttfamily. By default these are cmr, cmss and cmtt.

\bfdefault \mddefault

The series selected by \textbf, \bfseries, \textmd and \mdseries. By default these are bx and m. These values are suitable for the default families used. If other fonts are used as standard document fonts (for example, certain PostScript fonts) it might be necessary to adjust the value of \bfdefault to b since only a few such families have a ‘bold extended’ series. An alternative (taken for the fonts provided by psnfss) is to define silent substitutions from bx series to b series with special \DeclareFontShape declarations and the ssub size function, see Section 4.4.

\itdefault \sldefault \scdefault \updefault

The shapes selected by \textit, \itshape, \textsl, \slshape, \textsc,
\scshape, \textup and \upshape. By default these are it, sl, sc and n.

Note that there are no parameters for the size commands. These should be defined directly in class files, for example:

   \renewcommand{\normalsize}{\fontsize{10}{12}\selectfont}

More elaborate examples (setting additional parameters when the text size is changed) can be found in classes.dtx the source documentation for the classes

article, report, and book.

2.5 Special font declaration commands

\DeclareFixedFont {⟨cmd⟩} {⟨ENC ⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩} {⟨size⟩}

Declares command ⟨cmd⟩ to be a font switch which selects the font that is specified by the attributes ⟨ENC ⟩, ⟨family⟩, ⟨series⟩, ⟨shape⟩, and ⟨size⟩.

The font is selected without any adjustments to baselineskip and other surrounding conditions.

This example makes {\picturechar .} select a small dot very quickly:

   \DeclareFixedFont{\picturechar}{OT1}{cmr}{m}{n}{5}
\DeclareTextFontCommand {⟨cmd⟩} {⟨font-switches⟩}

Declares command ⟨cmd⟩ to be a font command with one argument. The current font attributes are locally modified by ⟨font-switches⟩ and then the argument of ⟨cmd⟩ is typeset in the resulting new font.

Commands defined by \DeclareTextFontCommand automatically take care of any necessary italic correction (on either side).

The following example shows how \textrm is defined by the kernel.

   \DeclareTextFontCommand{\textrm}{\rmfamily}

To define a command that always typeset its argument in the italic shape of the main document font you could declare:

   \DeclareTextFontCommand{\normalit}{\normalfont\itshape}

This declaration can be used to change the meaning of a command; if ⟨cmd⟩ is already defined, a log that it has been redefined is put in the transcript file.

\DeclareOldFontCommand {⟨cmd⟩} {⟨text-switch⟩} {⟨math-switch⟩}

Declares command ⟨cmd⟩ to be a font switch (i.e. used with the syntax {⟨cmd⟩...}) having the definition ⟨text-switch⟩ when used in text and the definition ⟨math-switch⟩ when used in a formula. Math alphabet commands, like \mathit, when used within ⟨math-switch⟩ should not have an argument. Their use in this argument causes their semantics to change so that they here act as a font switch, as required by the usage of the ⟨cmd⟩.

This declaration is useful for setting up commands like \rm to behave as they did in L^{A}TEX 2.09. We strongly urge you not to misuse this declaration to invent new font commands.

The following example defines \it to produce the italic shape of the main document font if used in text and to switch to the font that would normally be produced by the math alphabet \mathit if used in a formula.

   \DeclareOldFontCommand{\it}{\normalfont\itshape}{\mathit}

This declaration can be used to change the meaning of a command; if ⟨cmd⟩ is already defined, a log that it has been redefined is put in the transcript file.

3 Math fonts

This section describes the commands available to class and package writers for specifying math fonts and math commands.

3.1 Math font attributes

The selection of fonts within math mode is quite different to that of text fonts.

Some math fonts are selected explicitly by one-argument commands such as \mathsf{max} or \mathbf{vec}: these are called math alphabets. These math alphabet commands affect only the font used for letters and symbols of type \mathalpha (see Section 3.6); other symbols within the argument will be left unchanged. The predefined math alphabets are:

Alphabet  Description Example
\mathnormal  default  abcXY Z
\mathrm  roman  abcXYZ
\mathbf  bold roman abcXYZ
\mathsf sans serif abcXYZ
\mathit  text italic abcXYZ

                  \mathtt  typewriter  abcXYZ

\mathcal  calligraphic  X YZ

Other math fonts are selected implicitly by TEX for symbols, with commands such as \oplus (producing ⊕) or with straight characters like > or +. Fonts containing such math symbols are called math symbol fonts. The predefined math symbol fonts are:

Symbol font  Description  Example

           operators  symbols from \mathrm  [ + ]
            letters  symbols from \mathnormal << ⋆ >>

symbols  most L^{A}TEX symbols  ≤ ∗ ≥
largesymbols  large symbols  ∑ ∏ ∫

Some math fonts are both math alphabets and math symbol fonts, for example

\mathrm and operators are the same font, and \mathnormal and letters are

the same font.

Math fonts in L^{A}TEX have the same five attributes as text fonts: encoding, family, series, shape and size. However, there are no commands that allow the attributes to be individually changed. Instead, the conversion from math fonts to these five attributes is controlled by the math version. For example, the normal math version maps:

Math font  External font
Alphabets Symbol fonts  Attributes

             \mathnormal  letters       OML cmm  m it
                 \mathrm  operators     OT1 cmr  m  n
                \mathcal  symbols       OMS cmsy m  n
                          largesymbols  OMX cmex m  n
                 \mathbf                 OT1 cmr bx n
                 \mathsf                 OT1 cmss m  n
                 \mathit                 OT1 cmr  m it
                 \mathtt                 OT1 cmtt m  n

The bold math version is similar except that it contains bold fonts. The command \boldmath selects the bold math version.

Math versions can only be changed outside of math mode.

The two predefined math versions are:

normal the default math version
bold the bold math version

Packages may define new math alphabets, math symbol fonts, and math versions. This section describes the commands for writing such packages.

3.2 Selection commands

There are no commands for selecting symbol fonts. Instead, these are selected indirectly through symbol commands like \oplus. Section 3.6 explains how to define symbol commands.

\mathnormal{⟨math⟩}
\mathcal{⟨math⟩}
\mathrm{⟨math⟩}
\mathbf{⟨math⟩}
\mathsf{⟨math⟩}
\mathit{⟨math⟩}
\mathtt{⟨math⟩}

Each math alphabet is a command which can only be used inside math mode.

For example, $x + \mathsf{y} + \mathcal{Z}$ produces x + y + Z.
\mathversion{⟨version-name⟩}

This command selects a math version; it can only be used outside math mode.

For example, \boldmath is defined to be \mathversion{bold}.

3.3 Declaring math versions

\DeclareMathVersion {⟨version-name⟩}

Defines ⟨version-name⟩ to be a math version.

The newly declared version is initialised with the defaults for all symbol fonts and math alphabets declared so far (see the commands \DeclareSymbolFont

and \DeclareMathAlphabet).

If used on an already existing version, an information message is written to the transcript file and all previous \SetSymbolFont or \SetMathAlphabet declarations for this version are overwritten by the math alphabet and symbol font defaults, i.e. one ends up with a virgin math version.

Example:

   \DeclareMathVersion{normal}

3.4 Declaring math alphabets

\DeclareMathAlphabet {⟨math-alph⟩} {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

Defines ⟨math-alph⟩ to be a new math alphabet.

The arguments ⟨encoding⟩ ⟨family⟩ ⟨series⟩ ⟨shape⟩ are the default values for this math alphabet in all math versions; these can be reset later for a particular math version by a \SetMathAlphabet command. If ⟨shape⟩ is empty then the ⟨math-alph⟩ is declared to be invalid in all versions, unless it is set by a later

\SetMathAlphabet command.

Checks that ⟨math-alph⟩ can be used and that ⟨encoding⟩ is a valid encoding scheme.

In these examples, \foo is defined everywhere but \baz, by default, is defined nowhere.

   \DeclareMathAlphabet{\foo}{OT1}{cmtt}{m}{n}
   \DeclareMathAlphabet{\baz}{OT1}{}{}{}

\SetMathAlphabet {⟨math-alph⟩} {⟨version-name⟩} {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

Changes, or sets, the font for the math alphabet ⟨math-alph⟩ in math version ⟨version-name⟩ to ⟨encoding⟩⟨family⟩⟨series⟩⟨shape⟩.

Checks that ⟨math-alph⟩ is a math alphabet, ⟨version-name⟩ is a math version and ⟨encoding⟩ is a known encoding scheme.

This example defines \baz for the ‘normal’ math version only:

   \SetMathAlphabet\baz{normal}{OT1}{cmss}{m}{n}

Note that this declaration is not used for all math alphabets: Section 3.5 describes \DeclareSymbolFontAlphabet, which is used to set up math alphabets contained in fonts which have are declared as symbol fonts.

3.5 Declaring symbol fonts

\DeclareSymbolFont {⟨sym-font-name⟩} {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

Defines ⟨sym-font-name⟩ to be a new symbol font.

The arguments ⟨encoding⟩ ⟨family⟩ ⟨series⟩ ⟨shape⟩ are the default values for this symbol font in all math versions; these can be reset later for a particular math version by a \SetSymbolFont command.

Checks that ⟨sym-font-name⟩ can be used and that ⟨encoding⟩ is a declared encoding scheme.

For example, the following sets up the first four standard math symbol fonts:

   \DeclareSymbolFont{operators}{OT1}{cmr}{m}{n}
   \DeclareSymbolFont{letters}{OML}{cmm}{m}{it}
   \DeclareSymbolFont{symbols}{OMS}{cmsy}{m}{n}
   \DeclareSymbolFont{largesymbols}{OMX}{cmex}{m}{n}

\SetSymbolFont {⟨sym-font-name⟩} {⟨version name⟩} {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

Changes the symbol font ⟨sym-font-name⟩ for math version ⟨version name⟩ to ⟨encoding⟩ ⟨family⟩ ⟨series⟩ ⟨shape⟩.

Checks that ⟨sym-font-name⟩ is a symbol font, ⟨version name⟩ is a known math version and ⟨encoding⟩ is a declared encoding scheme.

For example, the following come from the set up of the ‘bold’ math version:

   \SetSymbolFont{operators}{bold}{OT1}{cmr}{bx}{n}
   \SetSymbolFont{letters}{bold}{OML}{cmm}{b}{it}
\DeclareSymbolFontAlphabet {⟨math-alph⟩} {⟨sym-font-name⟩}

Allows the previously declared symbol font ⟨sym-font-name⟩ to be also the math alphabet ⟨id⟩ (in all math versions).

This declaration should be used in preference to \DeclareMathAlphabet and \SetMathAlphabet when a math alphabet is the same as a symbol font; this is because it makes better use of the limited number (only 16) of TEX’s math groups.

Checks that ⟨math-alph⟩ can be defined and that ⟨sym-font-name⟩ is a symbol font.

Example:

   \DeclareSymbolFontAlphabet{\mathrm}{operators}
   \DeclareSymbolFontAlphabet{\mathcal}{symbols}

3.6 Declaring math symbols

\DeclareMathSymbol {⟨symbol⟩} {⟨type⟩} {⟨sym-font-name⟩} {⟨slot⟩}

The ⟨symbol⟩ can be either a single character such as ‘>’, or a macro name, such

as \sum.

Defines the ⟨symbol⟩ to be a math symbol of type ⟨type⟩ in slot ⟨slot⟩ of symbol font ⟨sym-font-name⟩. The ⟨type⟩ can be given as a number or as a command:

Type  Meaning  Example
0 or \mathord  Ordinary  α
1 or \mathop  Large operator  ∑
2 or \mathbin  Binary operation  ×
3 or \mathrel  Relation  ≤

            4 or \mathopen  Opening  ⟨
            5 or \mathclose  Closing  ⟩

6 or \mathpunct  Punctuation  ;
7 or \mathalpha Alphabet character  A

Only symbols of type \mathalpha will be affected by math alphabet commands: within the argument of a math alphabet command they will produce the character in slot ⟨slot⟩ of that math alphabet’s font. Symbols of other types will always produce the same symbol (within one math version).

\DeclareMathSymbol allows a macro ⟨symbol⟩ to be redefined only if it was previously defined to be a math symbol. It also checks that the ⟨sym-font-name⟩ is a declared symbol font.

Example:

   \DeclareMathSymbol{\alpha}{0}{letters}{"0B}
   \DeclareMathSymbol{\lessdot}{\mathbin}{AMSb}{"0C}
   \DeclareMathSymbol{\alphld}{\mathalpha}{AMSb}{"0C}

\DeclareMathDelimiter {⟨cmd⟩} {⟨type⟩} {⟨sym-font-name-1 ⟩} {⟨slot-1 ⟩} {⟨sym-font-name-2 ⟩} {⟨slot-2 ⟩}

Defines ⟨cmd⟩ to be a math delimiter where the small variant is in slot ⟨slot-1 ⟩ of symbol font ⟨sym-font-name-1 ⟩ and the large variant is in slot ⟨slot-2 ⟩ of symbol font ⟨sym-font-name-2 ⟩. Both symbol fonts must have been declared previously.

Checks that ⟨sym-font-name-i⟩ are both declared symbol fonts.

If TEX is not looking for a delimiter, ⟨cmd⟩ is treated just as if it had been defined with \DeclareMathSymbol using ⟨type⟩,⟨sym-font-name-1 ⟩ and ⟨slot-1 ⟩. In other words, if a command is defined as a delimiter then this automatically defines it as a math symbol.

Example:

   \DeclareMathDelimiter{\langle}{\mathopen}{symbols}{"68}
                                            {largesymbols}{"0A}

\DeclareMathAccent {⟨cmd⟩} {⟨type⟩} {⟨sym-font-name⟩} {⟨slot⟩}

Defines ⟨cmd⟩ to act as a math accent.

The accent character comes from slot ⟨slot⟩ in ⟨sym-font-name⟩. The ⟨type⟩ can be either \mathord or \mathalpha; in the latter case the accent character changes font when used in a math alphabet.

Example:

   \DeclareMathAccent{\acute}{\mathalpha}{operators}{"13}
   \DeclareMathAccent{\vec}{\mathord}{letters}{"7E}

\DeclareMathRadical {⟨cmd⟩} {⟨sym-font-name-1 ⟩} {⟨slot-1 ⟩} {⟨sym-font-name-2 ⟩} {⟨slot-2 ⟩}

Defines ⟨cmd⟩ to be a radical where the small variant is in slot ⟨slot-1 ⟩ of symbol font ⟨sym-font-name-1 ⟩ and the large variant is in slot ⟨slot-2 ⟩ of symbol font ⟨sym-font-name-2 ⟩. Both symbol fonts must have been declared previously.

Example (probably the only use for it!):

   \DeclareMathRadical{\sqrt}{symbols}{"70}{largesymbols}{"70}

3.7 Declaring math sizes

\DeclareMathSizes {⟨t-size⟩} {⟨mt-size⟩} {⟨s-size⟩} {⟨ss-size⟩}

Declares that ⟨mt-size⟩ is the (main) math text size, ⟨s-size⟩ is the ‘script’ size and ⟨ss-size⟩ the ‘scriptscript’ size to be used in math, when ⟨t-size⟩ is the current text size. For text sizes for which no such declaration is given the ‘script’ and ‘scriptscript’ size will be calculated and then fonts are loaded for the calculated sizes or the best approximation (this may result in a warning message).

Normally, ⟨t-size⟩ and ⟨mt-size⟩ will be identical; however, if, for example, PostScript text fonts are mixed with bit-map math fonts then you may not have available a ⟨mt-size⟩ for every ⟨t-size⟩.

Example:

   \DeclareMathSizes{13.82}{12.4}{10}{7}

4 Font installation

This section explains how L^{A}TEX’s font attributes are turned into TEX font specifications.

4.1 Font definition files

The description of how L^{A}TEX font attributes are turned into TEX fonts is usually kept in a font definition (.fd) file. The .fd file for family ⟨family⟩ in encoding ⟨encoding⟩ is called ⟨ENC ⟩⟨family⟩.fd, for example OT1cmr.fd for Computer Modern Roman or T1ptm.fd for Adobe Times.

Whenever L^{A}TEX encounters an encoding/family combination that it does not know (e.g. if the document designer says \fontfamily{ptm}\selectfont) then L^{A}TEX attempts to load the appropriate .fd file. “Not known” means: there was no \DeclareFontFamily declaration issued for this encoding/family combination.

The declarations in the .fd file are responsible for telling L^{A}TEX how to load fonts for that encoding/family combination. If the .fd file could not be found, a warning is issued and font substitutions are made.

4.2 Font definition file commands

The .fd files should contain only commands from this section. Note that those commands can also be used outside a .fd file: they can be put in package or class files, or even in the preamble of a document.

\ProvidesFile{⟨file-name⟩}[⟨release-info⟩]

The .fd file should announce itself with a \ProvidesFile command, as described in L^{A}TEX 2ε for Class and Package Writers. For example:

   \ProvidesFile{T1ptm.fd}[1994/06/01 Adobe Times font definitions]

\DeclareFontFamily {⟨encoding⟩} {⟨family⟩} {⟨loading-settings⟩}

Declares a font family ⟨family⟩ to be available in encoding scheme ⟨encoding⟩.

The ⟨loading-settings⟩ are executed immediately after loading any font with this encoding and family.

Checks that ⟨encoding⟩ was previously declared.

This example refers to the Computer Modern Typewriter font family in the Cork encoding:

   \DeclareFontFamily{T1}{cmtt}{\hyphenchar\font=-1}

Each .fd file should contain exactly one \DeclareFontFamily command, and it should be for the appropriate encoding/family combination.

\DeclareFontShape {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩} {⟨loading-info⟩} {⟨loading-settings⟩}

Declares a font shape combination; here ⟨loading-info⟩ contains the information that combines sizes with external fonts. The syntax is complex and is described in Section 4.3 below.

The ⟨loading-settings⟩ are executed after loading any font with this font shape. They are executed immediately after the ‘loading-settings’ which were declared by \DeclareFontFamily and so they can be used to overwrite the settings made at the family level.

Checks that the combination ⟨encoding⟩⟨family⟩ was previously declared via

\DeclareFontFamily.

Example:

   \DeclareFontShape{OT1}{cmr}{m}{sl}{%
             <5-8> sub * cmr/m/n
             <8> cmsl8
             <9> cmsl9
             <10> <10.95> cmsl10
             <12> <14.4> <17.28> <20.74> <24.88> cmsl12
             }{}

An .fd file can contain any number of \DeclareFontShape commands, which should be for the appropriate ⟨encoding⟩ and ⟨family⟩.

The font family declarations for the OT1-encoded fonts now all contain:

New feature 1996/06/01

  \hyphenchar\font=`\-

This enables the use of an alternative \hyphenchar in other encodings whilst maintaining the correct value for all fonts.

4.3 Font file loading information

The information which tells L^{A}TEX exactly which font (.tfm) files to load is contained in the ⟨loading-info⟩ part of a \DeclareFontShape declaration. This part consists of one or more ⟨fontshape-decl⟩s, each of which has the following form:

⟨fontshape-decl⟩ ::= ⟨size-infos⟩ ⟨font-info⟩
⟨size-infos⟩ ::= ⟨size-infos⟩ ⟨size-info⟩ | ⟨size-info⟩
⟨size-info⟩ ::= “<” ⟨number-or-range⟩ “>”
⟨font-info⟩ ::= [ ⟨size-function⟩ “*” ] [ “[” ⟨optarg⟩ “]” ] ⟨fontarg⟩

The ⟨number-or-range⟩ denotes the size or size-range for which this entry applies.

If it contains a hyphen it is a range: lower bound on the left (if missing, zero implied), upper bound on the right (if missing, ∞ implied). For ranges, the upper bound is not included in the range and the lower bound is.

Examples:

<10>  simple size 10pt only
<-8> range all sizes less than 8pt
<8-14.4> range  all sizes greater than or equal to 8pt

but less than 14.4pt

<14.4-> range  all sizes greater than or equal 14.4pt

If more than one ⟨size-info⟩ entry follows without any intervening ⟨font-info⟩, they all share the next ⟨font-info⟩.

The ⟨size-function⟩, if present, handles the use of ⟨font-info⟩. If not present, the ‘empty’ ⟨size-function⟩ is assumed.

All the ⟨size-info⟩s are inspected in the order in which they appear in the font shape declaration. If a ⟨size-info⟩ matches the requested size, its ⟨size-function⟩ is executed. If \external@font is non-empty afterwards this process stops, otherwise the next ⟨size-info⟩ is inspected. (See also \DeclareSizeFunction.)

If this process does not lead to a non-empty \external@font, L^{A}TEX tries the nearest simple size. If the entry contains only ranges an error is returned.

4.4 Size functions

L^{A}TEX provides the following size functions, whose ‘inputs’ are ⟨fontarg⟩ and ⟨optarg⟩ (when present).

‘’ (empty) Load the external font ⟨fontarg⟩ at the user-requested size. If ⟨optarg⟩ is present, it is used as the scale-factor.

s Like the empty function but without terminal warnings, only loggings.

gen Generates the external font from ⟨fontarg⟩ followed by the user-requested

     size, e.g. <8> <9> <10> gen * cmtt

sgen Like the ‘gen’ function but without terminal warnings, only loggings.

genb Generates the external font from ⟨fontarg⟩ followed by the user-requested

New feature 1995/12/01

size, using the conventions of the ‘ec’ fonts. e.g. <10.98> genb * dctt produces dctt1098.

sgenb Like the ‘genb’ function but without terminal warnings, only loggings.

New feature 1995/12/01

sub Tries to load a font from a different font shape declaration given by ⟨fontarg⟩ in the form ⟨family⟩/⟨series⟩/⟨shape⟩.

ssub Silent variant of ‘sub’, only loggings.

subf Like the empty function but issues a warning that it has to substitute the external font ⟨fontarg⟩ because the desired font shape was not available in the requested size.

ssubf Silent variant of ‘subf’, only loggings.

fixed Load font ⟨fontarg⟩ as is, disregarding the user-requested size. If present, ⟨optarg⟩ gives the “at … pt” size to be used.

sfixed Silent variant of ‘fixed’, only loggings.

Examples for the use of most of the above size functions can be found in the file cmfonts.fdd—the source for the standard .fd files describing the Computer Modern fonts by Donald Knuth.

\DeclareSizeFunction {⟨name⟩} {⟨code⟩}

Declares a size-function ⟨name⟩ for use in \DeclareFontShape commands. The interface is still under development but there should be no real need to a define new size functions.

The ⟨code⟩ is executed when the size or size-range in \DeclareFontShape matches the user-requested size.

The arguments of the size-function are automatically parsed and placed into \mandatory@arg and \optional@arg for use in ⟨code⟩. Also available, of course, is \f@size, which is the user-requested size.

To signal success ⟨code⟩ must define the command \external@font to contain the external name and any scaling options (if present) for the font to be loaded.

This example sets up the ‘empty’ size function (simplified):

   \DeclareSizeFunction{}
           {\edef\external@font{\mandatory@arg\space at\f@size}

5 Encodings

This section explains how to declare and use new font encodings and how to declare commands for use with particular encodings.

5.1 The fontenc package

Users can select new font encodings using the fontenc package. The fontenc package has options for encodings; the last option becomes the default encoding. For example, to use the OT2 (Washington University Cyrillic encoding) and T1 encodings, with T1 as the default, an author types:

   \usepackage[OT2,T1]{fontenc}

This package loads the encoding definition files (⟨ENC ⟩enc.def files) for each font encoding ⟨ENC ⟩ given as an option but not already declared; it also sets \encodingdefault to be the last encoding in the option list.

L^{A}TEX currently predefines the OT1 and T1 text encodings, and provides the files OT1enc.def and T1enc.def. Other encoding set-ups might be added to the distribution at a later stage.

Thus the example above loads the file OT2enc.def and sets \encodingdefault to T1.

Note: If you wish to use T1-encoded fonts other than the ‘cmr’ family then you may need to load the package (e.g. times) that selects the fonts before loading fontenc (this prevents the system from attempting to load any T1-encoded fonts from the ‘cmr’ family).

5.2 Encoding definition files

Warning: Some aspects of the contents of font definition files are still under development. Therefore, the current versions of the files OT1enc.def and T1enc.def are temporary versions and should not be used as models for producing further such files. For further information you should read the documen-

tation in ltoutenc.dtx.

The declarations in the encoding definition file are responsible for declaring the encoding and telling L^{A}TEX how to produce characters in this encoding.

The ⟨ENC ⟩enc.def files should contain only commands from this section. As with the font definition file commands, it is also possible (although normally not necessary) to use these declarations directly within a class or package file.

\ProvidesFile{⟨file-name⟩}[⟨release-info⟩]

The ⟨ENC ⟩enc.def file should announce itself with a \ProvidesFile command, described in L^{A}TEX 2ε for Class and Package Writers. For example:

   \ProvidesFile{OT2enc.def}
                [1994/06/01 Washington University Cyrillic encoding]

\DeclareFontEncoding {⟨encoding⟩} {⟨text-settings⟩} {⟨math-settings⟩}

Declares a new encoding scheme ⟨encoding⟩.

The ⟨text-settings⟩ are declarations which are executed every time \selectfont changes the encoding to be ⟨encoding⟩.

The ⟨math-settings⟩ are similar but are for math alphabets. They are executed whenever a math alphabet with this encoding is called.

Spaces within the arguments are ignored to avoid surplus spaces in the document. If a real space is necessary use \space.

Example:

   \DeclareFontEncoding{OT1}{}{}

Some author commands need to change their definition depending on which encoding is currently in use. For example, in the OT1 encoding, the letter ‘Æ’ is in slot "1D, whereas in the T1 encoding it is in slot "C6. So the definition of \AE has to change depending on whether the current encoding is OT1 or T1. The following commands allow this to happen.

\DeclareTextCommand {⟨cmd⟩} {⟨encoding⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩}

This command is like \newcommand, except that it defines a command which is specific to one encoding. For example, the definition of \aa in the OT1 encoding is:

  \DeclareTextCommand{\aa}{OT1}{{\accent23a}}

\DeclareTextCommands takes the same optional arguments as \newcommand.

The resulting command is robust, even if the code in ⟨definition⟩ is fragile.

It does not produce an error if the command has already been defined but logs the redefinition in the transcript file.

\ProvideTextCommand {⟨cmd⟩} {⟨encoding⟩} [⟨num⟩] [⟨default⟩] {⟨definition⟩} New feature 1994/12/01

This command is the same as \DeclareTextCommand, except that if ⟨cmd⟩ is already defined in encoding ⟨encoding⟩, then the definition is ignored.

\DeclareTextSymbol {⟨cmd⟩} {⟨encoding⟩} {⟨slot⟩}

This command defines a text symbol with slot ⟨slot⟩ in the encoding. For example, the definition of \ss in the OT1 encoding is:

   \DeclareTextSymbol{\ss}{OT1}{25}

It does not produce an error if the command has already been defined but logs the redefinition in the transcript file.

\DeclareTextAccent {⟨cmd⟩} {⟨encoding⟩} {⟨slot⟩}

This command declares a text accent, with the accent taken from slot ⟨slot⟩ in the encoding. For example, the definition of \" in the OT1 encoding is:

   \DeclareTextAccent{\"}{OT1}{127}

It does not produce an error if the command has already been defined but logs the redefinition in the transcript file.

\DeclareTextComposite {⟨cmd⟩} {⟨encoding⟩} {⟨letter⟩} {⟨slot⟩}

This command declares that the composite letter formed from applying ⟨cmd⟩ to ⟨letter⟩ is defined to be simply slot ⟨slot⟩ in the encoding. The ⟨letter⟩ should be a single letter (such as a) or a single command (such as \i).

For example, the definition of \'{a} in the T1 encoding could be declared like this:

   \DeclareTextComposite{\'}{T1}{a}{225}

The ⟨cmd⟩ should have been previously declared for this encoding, either with

using \DeclareTextAccent, or as a one-argument \DeclareTextCommand.
\DeclareTextCompositeCommand {⟨cmd⟩} {⟨encoding⟩} {⟨letter⟩} {⟨definition⟩}

New feature 1994/12/01

This is a more general form of \DeclareTextComposite, which allows for an arbitrary ⟨definition⟩, not just a ⟨slot⟩. The main use for this is to allow accents on i to act like accents on \i, for example:

   \DeclareTextCompositeCommand{\'}{OT1}{i}{\'\i}

It has the same restrictions as \DeclareTextComposite.

5.3 Default definitions

The commands in ⟨ENC ⟩enc.def files allow encoding-specific commands to be New description defined, but they do not allow commands to be used in other encodings. For 1994/12/01 example, the OMS encoding contains the symbol ‘§’, but we need to be able to use the command \S in any encoding, not just OMS.

To allow this, L^{A}TEX has commands for giving default definitions for commands, which are used when the command is not defined in the current encoding. For example, the default encoding for \S is OMS, and so in an encoding (such as OT1) which does not contain \S, the OMS version is selected. But in an encoding (such as T1) which does contain \S, the version for that encoding is used.

Note: These commands should not occur in ⟨ENC ⟩enc.def files, since these should only define commands for that encoding. They should instead be placed in packages.

\DeclareTextCommandDefault {⟨cmd⟩} {⟨definition⟩}

New feature 1994/12/01

This command allows an encoding-specific command to be given a default definition. For example, the default definition for \copyright is defined be be a circled ‘c’ with:

   \DeclareTextCommandDefault{\copyright}{\textcircled{c}}
\DeclareTextAccentDefault {⟨cmd⟩} {⟨encoding⟩}
\DeclareTextSymbolDefault {⟨cmd⟩} {⟨encoding⟩}

New feature 1994/12/01

These commands allow an encoding-specific command to be given a default encoding. For example, the default encoding for \" and \ae is set to be OT1 by:

   \DeclareTextAccentDefault{\"}{OT1}
   \DeclareTextSymbolDefault{\ae}{OT1}

Note that \DeclareTextAccentDefault can be used on any one-argument encoding-specific command, not just those defined with \DeclareTextAccent. Similarly, \DeclareTextSymbolDefault can be used on any encoding-specific command with no arguments, not just those defined with \DeclareTextSymbol.

For more examples of these definitions, see ltoutenc.dtx.

\ProvideTextCommandDefault {⟨cmd⟩} {⟨definition⟩}

New feature 1994/12/01

This command is the same as \DeclareTextCommandDefault, except that if the command already has a default definition, then the definition is ignored. This is useful to give ‘faked’ definitions of symbols which may be given ‘real’ definitions by other packages. For example, a package might give a fake definition of

\textonequarter by saying:
   \ProvideTextCommandDefault{\textonequarter}{$\m@th\frac14$}

5.4 Encoding defaults

\DeclareFontEncodingDefaults {⟨text-settings⟩} {⟨math-settings⟩}

Declares ⟨text-settings⟩ and ⟨math-settings⟩ for all encoding schemes. These are executed before the encoding scheme dependent ones are executed so that one can use the defaults for the major cases and overwrite them if necessary using

\DeclareFontEncoding.

If \relax is used as an argument, the current setting of this default is left unchanged.

This example is used by amsfonts.sty for accent positioning; it changes only the math settings:

   \DeclareFontEncodingDefaults{\relax}{\def\accentclass@{7}}

\DeclareFontSubstitution {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩}

Declares the default values for font substitution which will be used when a font with encoding ⟨encoding⟩ should be loaded but no font can be found with the current attributes.

These substitutions are local to the encoding scheme because the encoding scheme is never substituted! They are tried in the order ⟨shape⟩ then ⟨series⟩ and finally ⟨family⟩.

If no defaults are set up for an encoding, the values given by \DeclareErrorFont are used.

The font specification for ⟨encoding⟩⟨family⟩⟨series⟩⟨shape⟩ must have been de-

fined by \DeclareFontShape before the \begin{document} is reached.

Example:

   \DeclareFontSubstitution{T1}{cmr}{m}{n}

5.5 Case changing

\MakeUppercase {⟨text⟩}
\MakeLowercase {⟨text⟩}

TEX provides the two primitives \uppercase and \lowercase for changing the

New feature 1995/06/01

case of text. Unfortunately, these TEX primitives do not change the case of characters accessed by commands like \ae or \aa. To overcome this problem, L^{A}TEX provides these two commands.

In the long run, we would like to use all-caps fonts rather than any command like \MakeUppercase but this is not possible at the moment because such fonts do not exist.

For further details, see clsguide.tex.

In order that upper/lower-casing will work reasonably well, and in order to New description provide any correct hyphenation, L^{A}TEX 2ε must use, throughout a document, 1995/12/01 the same fixed table for changing case. The table used is designed for the font encoding T1; this works well with the standard TEX fonts for all Latin alphabets but will cause problems when using other alphabets.

6 Miscellanea

This section covers the remaining font commands in L^{A}TEX and some other issues.

6.1 Font substitution

\DeclareErrorFont {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩} {⟨size⟩}

Declares ⟨encoding⟩⟨family⟩⟨series⟩⟨shape⟩ to be the font shape used in cases where the standard substitution mechanism fails (i.e. would loop). For the standard mechanism see the command \DeclareFontSubstitution above.

The font specification for ⟨encoding⟩⟨family⟩⟨series⟩⟨shape⟩ must have been de-

fined by \DeclareFontShape before the \begin{document} is reached.

Example:

   \DeclareErrorFont{OT1}{cmr}{m}{n}{10}
\fontsubfuzz

This parameter is used to decide whether or not to produce a terminal warning if a font size substitution takes place. If the difference between the requested and the chosen size is less than \fontsubfuzz the warning is only written to the transcript file. The default value is 0.4pt. This can be redefined with

\renewcommand, for example:
   \renewcommand{\fontsubfuzz}{0pt}   % always warn

6.2 Preloading

\DeclarePreloadSizes {⟨encoding⟩} {⟨family⟩} {⟨series⟩} {⟨shape⟩} {⟨size-list⟩}

Specifies the fonts that should be preloaded by the format. These commands should be put in a preload.cfg file, which is read in when the L^{A}TEX format is being built. Read preload.dtx for more information on how to built such a configuration file.

Example:

   \DeclarePreloadSizes{OT1}{cmr}{m}{sl}{10,10.95,12}

6.3 Accented characters

Accented characters in L^{A}TEX can be produced using commands such as \"a etc. New description The precise effect of such commands depends on the font encoding being used. 1996/06/01 When using a font encoding that contains the accented characters as individual glyphs (such as the T1 encoding, in the case of \"a) words that contain such accented characters can be automatically hyphenated. For font encodings that do not contain the requested individual glyph (such as the OT1 encoding) such a command invokes typesetting instructions that produce the accented character as a combination of character glyphs and diacritical marks in the font. In most cases this involves a call to the TEX primitive \accent. Glyphs constructed as composites in this way inhibit hyphenation of the current word; this is one reason why the T1 encoding is preferable to the original TEX font encoding OT1.

It is important to understand that commands like \"a in L^{A}TEX 2ε represent just a name for a single glyph (in this case ‘umlaut a’) and contain no information about how to typeset that glyph—thus it does not mean ‘put two dots on top of the character a’. The decision as to what typesetting routine to use will depend on the encoding of the current font and so this decision is taken at the last minute. Indeed, it is possible that the same input will be typeset in more than one way in the same document; for example, text in section headings may also appear in table of contents and in running heads; and each of these may use a font with a different encoding.

For this reason the notation \"a is not equivalent to:

  \newcommand \chara {a}     \"\chara

In the latter case, L^{A}TEX does not expand the macro \chara but simply compares the notation (the string \"\chara) to its list of known composite notations in the current encoding; when it fails to find \"\chara it does the best it can and invokes the typesetting instructions that put the umlaut accent on top of the expansion of \chara. Thus, even if the font actually contains ‘ä’ as an individual glyph, it will not be used.

The low-level accent commands in L^{A}TEX are defined in such a way that it is possible to combine a diacritical mark from one font with a glyph from another font; for example, \"\textparagraph will produce ¨¶. The umlaut here is taken from the OT1 encoded font cmr10 whilst the paragraph sign is from the OMS encoded font cmsy10. (This example may be typographically silly but better ones would involve font encodings like OT2 (Cyrillic) that might not be available at every site.)

There are, however, restrictions on the font-changing commands that will work within the argument to such an accent command. These are TEXnical in the sense that they follow from the way that TEX’s \accent primitive works, allowing only a special class of commands between the accent and the accented character.

The following are examples of commands that will not work correctly as the accent will appear above a space: the font commands with text arguments (\textbf{...} and friends); all the font size declarations (\fontsize and \Large, etc.); \usefont and declarations that depend on it, such as

\normalfont; box commands (e.g. \mbox{...}).

The lower-level font declarations that set the attributes family, series and shape (such as \fontshape{sl}\selectfont) will produce correct typesetting, as will the default declarations such as \bfseries.

6.4 Naming conventions

• Math alphabet commands all start with \math...: examples are \mathbf,

     \mathcal, etc.

• The text font changing commands with arguments all start with \text...: e.g. \textbf and \textrm. The exception to this is \emph, since it occurs very commonly in author documents and so deserves a shorter name.

• Names for encoding schemes are strings of up to three letters, all upper case. The L^{A}TEX3 project reserves the use of encodings starting with T (standard 256-long text encodings), M (standard 256-long math encodings), S (standard 256-long symbol encodings), OT (standard 128-long text encodings) and OM (standard 128-long math encodings). Please do not use the above starting letters for non-portable encodings. If new standard encoding emerge then we shall add them in a later release of L^{A}TEX.

Encoding schemes which are local to a site should start with L.

• Font family names should contain up to five lower case letters. Where possible, these should conform to the Filenames for fonts font naming scheme.

• Font series names should contain up to four lower case letters.

• Font shapes should contain up to two letters lower case.

• Names for symbol fonts are built from lower and upper case letters with no restriction.

Whenever possible, you should use the series and shape names suggested in The L^{A}TEX Companion since this will make it easier to combine new fonts with existing fonts.

Where possible, text symbols should be named as \text followed by the Adobe New description glyph name: for example \textonequarter or \textsterling. Similarly, math 1994/12/01 symbols should be named as \math followed by the glyph name, for example \mathonequarter or \mathsterling. Commands which can be used in text or math can then be defined using \ifmmode, for example:

\DeclareRobustCommand{\pounds}{%
 \ifmmode \mathsterling \else \textsterling \fi
}

Note that commands defined in this way must be robust, in case they get put into a section title or other moving argument.

7 If you need to know more …

The tracefnt package provides for tracing the actions concerned with loading, New description substituting and using fonts. The package accepts the following options: 1996/06/01 errorshow Write all information about font changes, etc. but only to the tran-

script file unless an error occurs. This means that information about font substitution will not be shown on the terminal.

warningshow Show all font warnings on the terminal. This setting corresponds to the default behaviour when this tracefnt package is not used!

infoshow Show all font warnings and all font info messages (that are normally only written to the transcript file) also on the terminal. This is the default when this tracefnt package is loaded.

debugshow In addition to what is shown by infoshow, show also changes of math fonts (as far as possible): beware, this option can produce a large amount of output.

loading Show the names of external font files when they are loaded. This option shows only ‘newly loaded’ fonts, not those already preloaded in the format or the class file before this tracefnt package becomes active.

pausing Turn all font warnings into errors so that L^{A}TEX will stop.

Warning: the actions of this package can change the layout of a document and even, in rare cases produce clearly wrong output, so it should not be used in the final formatting of ‘real documents’.

References

[1] Michel Goossens, Frank Mittelbach and Alexander Samarin. The L^{A}TEX Companion. Addison-Wesley, Reading, Massachusetts, 1994.

[2] Donald E. Knuth. Typesetting concrete mathematics. TUGboat, 10(1):31– 36, April 1989.

[3] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1994.